### v2

@@ 5 Proof: existence and uniqueness
Furthermore, the stability result in Theorem 4.6 implies that
$\|Y^{n}-Y^{m}\|_{\mathbb{S}^{2p}}+\|Z^{n}-Z^{m}\|_{\mathbb{H}^{p}}\ \leq\ C_{p}\;\|\xi^{n}-\xi^{m}\|_{L^{2p}}$
where $C_{p}$ depends only on $(T,K_{z},K_{\tilde{z}},K_{0})$, thus is independent of $n$ and $m$.
Pick $(\xi^{n})_{n\in\mathbb{N}}$ is Cauchy in $L^{2p}$, it follows that $(Y^{n},Z^{n})_{n\in\mathbb{N}}$ is also Cauchy in $\mathbb{S}^{2p}\times\mathbb{H}^{p}$. As each $(Y^{n},Z^{n})$ lies in the Banach space $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$, we conclude the existence of a limit $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$. Passing to the limit in (5.28), we verify that the pair $(Y,Z)$ indeed solves the original mean-field BSDE (2.6), establishing existence. Uniqueness then follows directly from the stability result in Theorem 4.6, which concludes the proof. ∎
6 Application: Utility Maximization
@@
In this section, we provide an extension of the foundational work of hu2005utility to the mean-field case. Specifically, we explore how a trader’s decision-making process evolves in a financial market influenced by collective behavior. Unlike previous studies that assume fixed constraints, we consider a dynamic setting where a trader’s strategy is shaped not only by individual preferences but also by the actions of others. This interaction leads to a more flexible and realistic model, where part of the trading constraints are given by market conditions rather than being imposed arbitrarily.
@@
This section is structured as follows: In Section 6.1, we introduce the utility maximization problem and the associated mathematical framework. We then extend the given constraint set to incorporate a dependence on the stochastic process. It is important to note that, at this stage, we do not make any assumptions about this process beyond considering it as a general stochastic process. Next, we establish the connection between the utility maximization problem and the BSDE (6.30) with a thoughtfully chosen driver. We will subsequently demonstrate that the process defined earlier for the constraint set is precisely the solution $Z$ in the BSDE (6.30). With the complete framework in place, we offer a financial interpretation of the components, which is discussed in Remarks 6.4.
To connect this application to our earlier analysis about BSDE, we generalize the separately quadratic driver from BSDE (2.6) to a fully coupled quadratic driver in BSDE (6.32) and establish the existence and uniqueness of solutions in Proposition 6.7 when the terminal condition is small enough. To do this, we adopt a localization method in 6.6, then we extend it to the global case. Then, after we established the existence and uniqueness for BSDE (6.32), we then connect this back to the utility maximization problem by presenting the final result in Theorem 6.9. We first derive a key estimate in Lemma 6.3. This estimate is then used in Theorem 6.9 to verify that the driver of BSDE (6.30) satisfies conditions $(\rm H_{0})$} and $(\rm H_{q}^{\prime})$}, ensuring the well-posedness of BSDE (6.32). Leveraging these lemmas, we conclude with the main result in Theorem 6.9, which is supported by the proof of Theorem 7 in hu2005utility.
@@ 6.1 Financial market model setup
where $W$ is $n$-dimensional Brownian motion, $b^{i}$ is a real-valued predictable, bounded stochastic process, representing the drift, and $\sigma^{i}$ is an $\mathbb{R}^{n}$-valued predictable bounded stochastic process, representing the volatility. The volatility matrix $\sigma_{t}=(\sigma_{t}^{1},\dots,\sigma_{t}^{d})^{\top}$ is assumed to have full rank $d$. And we assume that $\sigma_{t}\sigma_{t}^{\top}$ is uniformly elliptic. Thus market price of risk $\theta_{t}$, also referred to as the risk premium, is defined by:
@@
For the investors other than the representative, we assume the risk aversion parameter $\alpha>0$ is identical. The model for the risky assets is of the same type but with $b,\sigma,W$ being independent and identically distributed to the representative’s model. In addition, we assume the terminal liabilities are independent and have distribution $F$.
@@ 6.2 Constraint set and aggregated market signal
To ensure that no arbitrage opportunities exist, the definition of admissible trading strategies imposes constraints on the set of allowable positions. Below, we outline the constraint set in Definition 6.1, which accounts for market-wide influences. To incorporate the influence of market-wide behavior, for $t\in[0,T]$ and $\omega\in\Omega$, we define two stochastic processes that satisfy the necessary integrability conditions such that
@@
where $Z_{t}(\omega)\in\mathbb{R}^{1\times n}$ and $\hat{Z_{t}}(\omega)\in\mathbb{R}^{1\times d}$. These two stochastic processes are essentially equivalent because the matrix $\sigma$ has full rank. Later on, in Remark 6.4, we will see $\hat{Z}_{t}\sigma_{t}+\frac{\theta_{t}}{\alpha}=Z_{t}+\frac{\theta_{t}}{\alpha}$ yields the optimal strategy at time $t\in[0,T]$ if no constants are imposed. Thus $Z$ can be interpreted as the trading strategy adjusted by $\theta/\alpha$. Later, we will show that this $Z$ will precisely correspond to the martingale density process of BSDE (6.30) as well.
We now present the following expressions for the constraint set under different settings. Specifically, we denote the constraint without market influence by $\hat{D}\subseteq\mathbb{R}^{d}$, where $\hat{D}$ is a closed set. When market influence is present, the strategy $\pi_{t}(\omega)$ is constrained to a new closed set.
@@
With the help of the previous definitions, for a fixed $t\in[0,T]$, we start with a set $\hat{D}\subset\mathbb{R}^{d}$, which represents certain constraints, such as risk limits or position constraints. Upon incorporating information from the market, we could apply scaling and shifting to this set, which now reflects both the original constraints and the market’s influence. This results in a new set that is much more flexible and realistic, capturing the dynamic behavior of the market. Therefore, we denote this updated set as $\hat{C}(\mathbb{E}[\hat{Z}_{t}])$, emphasizing its dependence on the aggregated market signal $\mathbb{E}[\hat{Z_{t}}$] at time $t\in[0,T]$.
@@
Note $Z_{t}=\hat{Z_{t}}\sigma_{t}$ and $\mathbb{E}[Z_{t}]=\mathbb{E}[\hat{Z_{t}}]\sigma_{t}$.
@@
$\sup_{p\in\mathcal{A}}\mathbb{E}\left[-\exp\left(-\alpha\left(x+\int_{0}^{T}p_{t}(\mathop{}\!dW_{t}+\theta_{t}\mathop{}\!dt)-F\right)\right)\right]=\mathbb{E}\left[-\exp(-\alpha(X_{T}^{p^{*}}-F))\right]$
@@
where $(Y,Z)$ is a solution of the BSDE
$Y_{t}=F+\int_{t}^{T}f(s,Z_{s},\mathbb{E}[Z_{s}])\,ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s},\quad t\in[0,T],.$ (6.30)
Here, the driver $f$ to be chosen such that $V^{(p)}(\cdot,x)$ is a supermartingale for all $p\in\mathcal{A}$, and there exists $p^{*}\in\mathcal{A}$ such that $V^{(p^{*})}(\cdot,x)$ is a true martingale. If we are able to choose such $f$, which will be shown later, we have
$\mathbb{E}V^{(p)}(T,x)\leq V^{(p)}(0,x)=V^{(p^{*})}(0,x)=\mathbb{E}V^{(p^{*})}(T,x).$
@@
Earlier, we described the process $(Z_{t})_{0\leq t\leq T}$ simply as a stochastic process. However, in this context, we now define it as the martingale density process $(Z_{t})_{0\leq t\leq T}$ for the BSDE in equation (6.30). Consequently, as the constraint set depends on $\mathbb{E}[Z_{t}]$, the driver should has dependence on $\mathbb{E}[Z_{t}]$ as well: see (6.31). To verify this, let us specify the driver $f$ for the construction of $\left\{V^{(p)}(\cdot,x):p\in\mathcal{A}\right\}$. By Itô’s formula,
@@
is a true martingale as $\alpha(Z-p)$ is square integrable. We require $\tilde{A}^{p}$ to be a non-increasing process such that $v(t,p_{t},Z_{t})\geq 0$ for all $p\in\mathcal{A}$, and there exists a $p^{*}\in\mathcal{A}$ such that $v(t,p^{*}_{t},Z_{t})=0$. Therefore, if we choose
@@
Thus, a family of processes $\left\{V^{(p)}(\cdot,x):p\in\mathcal{A}\right\}$ becomes a colletion of supermartingales where, for $p_{t}^{*}\in\Pi_{C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)}\left(Z_{t}+\frac{\theta_{t}}{\alpha}\right)$, we have $v(\cdot,p^{*},Z)=0$ for every $t\in[0,T]$. Therefore, $V^{(p^{*})}(\cdot,x)$ becomes a true martingale.
We have now completed the construction, and it is clear how the utility maximization problem in our setting connects to a type of mean-field type BSDEs as in hibon2017MFquadratic. In hu2005utility, the optimization problem was solved under the assumption that the random endowment $F$ is bounded and that the trading strategy $p$ takes values in the admissible set $\mathcal{A}$. However, this set differs from ours because, in their case, the constraint set for trading strategies is simply closed in $\mathbb{R}^{d}$. In contrast, we introduce a particular constraint set where the trader’s behavior depends on the market (i.e., the behavior of other traders) in a mean-field sense. Specifically, we expand the driver to include the expectation of the control process $Z$. Thus, we define a constraint set that depends on time, reflecting the market’s influence at each moment. This approach introduces greater flexibility and a more realistic model. To establish the connection between the control process $Z$ in the BSDE (6.30) and the trading strategy $p\in\mathcal{A}$, we provide the following remarks, which bridge the mathematical framework with financial interpretations.
@@
Thus, the process $Z_{t}=p_{t}^{*}-\frac{\theta_{t}}{\alpha}$ can be interpreted as the unconstrained trading strategy adjusted by $\frac{\theta_{t}}{\alpha}$ at time $t\in[0,T]$. In general situation, the constraint set $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)$ depends on the average of the adjusted trading strategies among the market participants. The convergence is guaranteed by the homogeneity and exchangeability of market participants.

### v3

@@ 5 Proof: existence and uniqueness
Moreover, since $\phi^{n}(\omega)$ converges to $\phi(\omega)$ uniformly in $\omega$, we have $\left\|\xi^{n}-\xi\right\|_{L^{\infty}}\to 0$. Then, by Theorem 4.7, we obtain the convergence of $(Y^{n},Z^{n})$ to $(Y,Z)$ in $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{BMO}$. The uniqueness follows from Theorem 4.7 again. ∎
Now let us generalize above theorem to the version we stated in Section 2.
Theorem 5.5.
Assume that there is a constant $C$ such that $b:[0,T]\times\mathbb{R}^{k}\to\mathbb{R}^{k}$ and $\sigma:[0,T]\to\mathbb{R}^{k\times n}$ are measurable functions with
  • $|b(t,x)-b(t,x^{\prime})|\leq C|x-x^{\prime}|,|b(t,0)|\leq C$ for every $(t,x)\in[0,T]\times\mathbb{R}^{k}$, and
  • $C^{-1}|v|^{2}\leq v^{\intercal}(\sigma\sigma^{\intercal})(t)v\leq C|v|^{2}$ for all $v\in\mathbb{R}^{k}$ and $t\in[0,T]$.
Then, let $X$ be the strong solution of
$dX_{t}=b(t,X_{t})dt+\sigma(t)dW_{t};\qquad X_{0}=x_{0}.$
Then, we can replace $\phi(W_{[0,T]})$ to $\phi(X_{[0,T]})$ Theorem 5.4, and the statement still holds true. In addition, if $b(t,x)$ is uniformly bounded, we can replace $\phi((W_{t_{i}})_{i=1,2,\cdots,n}))$ to $\phi((X_{t_{i}})_{i=1,2,\cdots,n}))$ in Theorem 5.3 as well.
Proof.
First, let us consider Theorem 5.3. Without loss of generality, let us assume $k=1$: otherwise, we can use a dynamic change of coordinate to diagonalize $\sigma\sigma^{\intercal}(t)$ and treat each coordinate separately. First of all, note that $D_{t}X_{s}$ is bounded due to Lemma 4.2 of Cheridito and Nam (2014). Therefore, in the proof of Theorem 5.3,
$\xi^{m}:=\phi^{m}(X_{t_{1}},X_{t_{2}},\cdots,X_{t_{n}})$
has uniformly bounded Malliavin derivatives, which gives the corresponding solution $(Y^{m},Z^{m})$. If we show that $\int_{0}^{T}|H_{t}^{m}|^{2}ds<\infty$ for $H^{m}$ defined by the martingale representation of $\xi^{m}$, then Theorem 4.6 gives the existence and uniqueness of solution. Note that $H^{m}$ does not change under equivalent measure change. Let us define $\mathbb{Q}$ by
$\left.\frac{d\mathbb{Q}}{d\mathbb{P}}\right|_{t}=\mathcal{E}_{t}\left(\int_{0}^{\cdot}(\sigma^{-1}b)(t,X_{t})^{\intercal}dW_{t}\right).$
Then we can apply Girsanov’s theorem to conclude that $\mathbb{Q}$ is equivalent to $\mathbb{P}$ and that
$B_{t}=W_{t}+\int_{0}^{t}\left(\sigma^{-1}b\right)(s,X_{s})ds$
is $\mathbb{Q}$-Brownian motion. We have $dX_{t}=\sigma(t)dB_{t}$ and, under time change, $X_{t}=\tilde{B}_{\int_{0}^{t}|\sigma(s)|^{2}ds}$ for some Brownian motion $\tilde{B}$. Since $|\sigma|^{2}>C^{-1}>0$, there exist deterministic times $s_{i}$ such that $X_{t_{i}}=\tilde{B}_{s_{i}}$. Since
$\xi^{m}=\phi^{m}(X_{t_{1}},X_{t_{2}},\cdots,X_{t_{n}})=\phi^{m}(\tilde{B}_{s_{1}},\tilde{B}_{s_{2}},\cdots,\tilde{B}_{s_{n}})$
with $\int_{0}^{s_{n}}|\sigma(s)|^{2}ds=T$. This implies $s_{n}<\infty$ because $|\sigma|^{2}$ is bounded from below by $C^{-1}$. Using the Lemma 5.2, we obtain the uniform bound on $H^{m}$.
On the other hand, let us consider Theorem 5.4 where $k$ can be an arbitrary natural number. Note that we only need to prove that, for Lipschitz $\phi:C([0,T];\mathbb{R})\to\mathbb{R}$, the Malliavin derivative of $\phi(X_{[0,T]})$ is bounded: if this statement is true, $\phi^{n}$ in the proof of Theorem 5.4 has bounded Malliavin derivatives, so we can use the same argument to prove the existence and uniqueness.
Note that, if $\phi$ is Lipschitz, for linear interpolator $l^{m}:\mathbb{R}^{m}\to C([0,T];\mathbb{R})$ and $t_{i}=\frac{iT}{m}$, we have
$\mathbb{E}\left[\left|\phi(X_{[0,T]})-(\phi\circ l^{m})(X_{t_{1}},X_{t_{2}},\cdots,X_{t_{m}})\right|^{p}\right]$ $\leq C^{2}\mathbb{E}\left[\max_{i=1,2,\cdots,m}\sup_{t\in(t_{i},t_{i+1}]}|X_{t}-X_{t_{i}}|^{p}\right],$
for $p>2$. Note that we have
$\sup_{t\in(t_{i},t_{i+1}]}|X_{t}-X_{t_{i}}|^{p}$ $\leq\sup_{t\in(t_{i},t_{i+1}]}\left|\int_{t_{i}}^{t}b(s,X_{s})ds+\int_{t_{i}}^{t}\sigma(s)dW_{s}\right|^{p}$
$\leq 2^{p-1}\sup_{t\in(t_{i},t_{i+1}]}\left(\left|\int_{t_{i}}^{t}b(s,X_{s})ds\right|^{p}+\left|\int_{t_{i}}^{t}\sigma(s)dW_{s}\right|^{p}\right)$
$\leq 2^{p-1}\left((t_{i+1}-t_{i})^{p-1}\int_{t_{i}}^{t_{i+1}}|b(s,X_{s})|^{p}ds+\sup_{t\in(t_{i},t_{i+1}]}\left|\int_{t_{i}}^{t}\sigma(s)dW_{s}\right|^{p}\right)$
$\leq 2^{p-1}\left((2T)^{p-1}m^{-(p-1)}C^{p}\int_{0}^{T}(1+|X_{s}|^{2})ds+\sup_{t\in(t_{i},t_{i+1}]}\left|\int_{t_{i}}^{t}\sigma(s)dW_{s}\right|^{p}\right).$
It is easy to see that, as $m\to\infty$,
$\mathbb{E}\max_{i=1,2,\cdots,m}m^{-(p-1)}\int_{0}^{T}(1+|X_{s}|^{2})ds\leq\mathbb{E}m^{-(p-2)}\int_{0}^{T}(1+|X_{s}|^{2})ds\to 0$
because $X\in\mathbb{H}^{2}$ from our assumption. On the other hand
$\mathbb{E}\max_{i=1,2,\cdots,m}\sup_{t\in(t_{i},t_{i+1}]}\left|\int_{t_{i}}^{t}\sigma(s)dW_{s}\right|^{p}$ $\leq\sum_{i=1}^{m}\mathbb{E}\sup_{t\in(t_{i},t_{i+1}]}\left|\int_{t_{i}}^{t}\sigma(s)dW_{s}\right|^{p}\leq C_{p}\sum_{i=1}^{m}\mathbb{E}\left|\int_{t_{i}}^{t_{i+1}}|\sigma(s)|^{2}ds\right|^{p/2}$
$\leq C_{p}\sum_{i=1}^{m}(t_{i+1}-t_{i})^{p/2-1}\int_{t_{i}}^{t_{i+1}}|\sigma(s)|^{p}ds$
$\leq C_{p}\left(\frac{T}{m}\right)^{p/2-1}\int_{0}^{T}|\sigma(s)|^{p}ds\to 0,$
as $m\to\infty$. Here, $C_{p}$ is the constant from the Burkholder-Davis-Gundy inequality. As a result, $(\phi\circ l^{m})(X_{t_{1}},X_{t_{2}},\cdots,X_{t_{m}})\to\phi(X_{[0,T]})$ in $L^{p}$. In addition, by (Nualart, 2006, Proposition 1.2.4) and the boundedness of $D_{t}X_{s}$, we know that $(\phi\circ l^{m})(X_{t_{1}},X_{t_{2}},\cdots,X_{t_{m}})$ has bounded Malliavin derivative. Therefore, by (Nualart, 2006, Lemma 1.2.3), we conclude that $\phi(X_{[0,T]})$ has bounded Malliavin derivative whenever $\phi$ is Lipschitz. ∎
6 Multidimensional MFBSDE with Small Terminal Condition and Utility Maximization
@@
In this section, we provide an extension of the foundational work of Hu et al. (2005) to the mean-field case. Specifically, we explore how a trader’s decision-making process evolves in a financial market influenced by collective behavior. Unlike previous studies that assume fixed constraints, we consider a dynamic setting where a trader’s strategy is shaped not only by individual preferences but also by the actions of others. This interaction leads to a more flexible and realistic model, in which part of the trading constraints are determined by market conditions rather than imposed arbitrarily.
@@
This section is structured as follows. In Section 6.1, we introduce the utility maximization problem and the associated mathematical framework. We then extend the constraint set to incorporate dependence on a stochastic process. At this stage, we do not impose any assumptions on this process beyond considering it as a general stochastic process. Next, we establish the connection between the utility maximization problem and the BSDE (6.31) with a suitably chosen driver. We then show that the process used to define the constraint set is precisely the solution component $Z$ in BSDE (6.31). With the complete framework in place, we offer a financial interpretation of the components, which is discussed in Remarks 6.4.
To relate this application to our theoretical results, we move beyond the structural assumptions on the terminal condition used in Sections 45. The driver in (6.32) involves the distance to a constraint set that depends on the aggregated signal $\mathbb{E}[Z]$, and its quadratic growth is generally fully coupled in $(Z,\mathbb{E}[Z])$. Consequently, we fit (6.31) into the unified frameworks used earlier via a different representation. We introduce the fully coupled BSDE (6.33) and prove a dedicated well-posedness result under a small terminal oscillation assumption.
More precisely, we establish local existence and uniqueness for the centered BSDE (6.34) on short time intervals in Lemma 6.6 via a contraction argument, and then obtain a global solution on $[0,T]$ by a concatenation procedure in Proposition 6.7. The smallness requirement is imposed on the centered terminal condition $\eta:=\xi-\mathbb{E}[\xi]$, which is natural in the utility maximization setting where the terminal liability need not be small but may be close to its mean.
With this well-posedness result in hand, we return to the utility maximization problem. We first derive an elementary growth estimate for the (volatility-scaled) constraint set in Lemma 6.3. This estimate is used in Theorem 6.9 to verify that the driver $f$ in (6.32) satisfies $(\mathrm{H}_{0})$ and falls within the fully coupled quadratic class covered by Proposition 6.7. Finally, we conclude that the BSDE (6.31) admits a unique solution, and we identify the value function and an optimal strategy by adapting the verification argument in Theorem 7 of Hu et al. (2005) to the present mean-field, market-adaptive constraint setting.
@@ 6.1 Financial market model setup
where $W$ is an $n$-dimensional Brownian motion, $b^{i}$ is a real-valued predictable, bounded stochastic process representing the drift, and $\sigma^{i}$ is an $\mathbb{R}^{n}$-valued predictable, bounded stochastic process representing the volatility. The volatility matrix $\sigma_{t}=(\sigma_{t}^{1},\dots,\sigma_{t}^{d})^{\top}$ is assumed to have full rank $d$, and we assume that $\sigma_{t}\sigma_{t}^{\top}$ is uniformly elliptic. Thus, the market price of risk $\theta_{t}$, also referred to as the risk premium, is defined by
@@
For investors other than the representative, we assume that the risk aversion parameter $\alpha>0$ is identical. The model for the risky assets is of the same type, but with $b$, $\sigma$ and $W$ being independent and identically distributed relative to the representative’s model. In addition, we assume that the terminal liabilities are independent and have distribution $F$.
@@ 6.2 Constraint set and aggregated market signal
To ensure that no arbitrage opportunities exist, the definition of admissible trading strategies imposes constraints on the set of allowable positions. Below, we outline the constraint set in Definition 6.1, which accounts for market-wide influences. To incorporate the influence of market-wide behavior, for $t\in[0,T]$ and $\omega\in\Omega$, we define two stochastic processes satisfying the necessary integrability conditions such that
@@
where $Z_{t}(\omega)\in\mathbb{R}^{1\times n}$ and $\hat{Z_{t}}(\omega)\in\mathbb{R}^{1\times d}$. These two processes are essentially equivalent because the matrix $\sigma$ has full rank. Later, in Remark 6.4, we will see that $\hat{Z}_{t}\sigma_{t}+\frac{\theta_{t}}{\alpha}=Z_{t}+\frac{\theta_{t}}{\alpha}$ yields the optimal strategy at time $t\in[0,T]$ if no constraints are imposed. Thus, $Z$ can be interpreted as the trading strategy adjusted by $\theta/\alpha$. Later, we will show that this $Z$ corresponds precisely to the martingale density process of BSDE (6.31).
We now present expressions for the constraint set under different settings. Specifically, we denote the constraint without market influence by $\hat{D}\subseteq\mathbb{R}^{d}$, where $\hat{D}$ is a closed set. When market influence is present, the strategy $\pi_{t}(\omega)$ is constrained to a new closed set.
@@
With the help of the previous definitions, for a fixed $t\in[0,T]$, we start with a set $\hat{D}\subset\mathbb{R}^{d}$, which represents certain constraints such as risk limits or position constraints. After incorporating information from the market, we apply scaling and shifting to this set so that it reflects both the original constraints and the market’s influence. This results in a new set that is more flexible and realistic, capturing the dynamic behavior of the market. We denote this updated set by $\hat{C}(\mathbb{E}[\hat{Z}_{t}])$, emphasizing its dependence on the aggregated market signal $\mathbb{E}[\hat{Z}_{t}]$ at time $t\in[0,T]$.
@@
Note that $Z_{t}=\hat{Z}_{t}\,\sigma_{t}$ and $\mathbb{E}[Z_{t}]=\mathbb{E}[\hat{Z}_{t}]\,\sigma_{t}$.
@@
$\sup_{p\in\mathcal{A}}\mathbb{E}\left[-\exp\left(-\alpha\left(x+\int_{0}^{T}p_{t}(\mathop{}\!dW_{t}+\theta_{t}\mathop{}\!dt)-F\right)\right)\right]=\mathbb{E}\left[-\exp(-\alpha(X_{T}^{p^{*}}-F))\right],$
@@
where $(Y,Z)$ is a solution of the MFBSDE
$Y_{t}=F+\int_{t}^{T}f(s,Z_{s},\mathbb{E}[Z_{s}])\,ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s},\quad t\in[0,T].$ (6.31)
Here, the driver $f$ will be chosen so that $V^{(p)}(\cdot,x)$ is a supermartingale for all $p\in\mathcal{A}$, and there exists $p^{*}\in\mathcal{A}$ such that $V^{(p^{*})}(\cdot,x)$ is a true martingale. If we can choose such an $f$ (as we show below), then
$\mathbb{E}V^{(p)}(T,x)\leq V^{(p)}(0,x)=V^{(p^{*})}(0,x)=\mathbb{E}V^{(p^{*})}(T,x),$
@@
Earlier, we described the process $(Z_{t})_{0\leq t\leq T}$ simply as a stochastic process. However, in this context, we now define it as the martingale density process $(Z_{t})_{0\leq t\leq T}$ for the BSDE in (6.31). Consequently, since the constraint set depends on $\mathbb{E}[Z_{t}]$, the driver should also depend on $\mathbb{E}[Z_{t}]$; see (6.32). To verify this, we now specify the driver $f$ for the construction of $\left\{V^{(p)}(\cdot,x):p\in\mathcal{A}\right\}$. By Itô’s formula,
@@
is a true martingale as $\alpha(Z-p)$ is square integrable. We require $\tilde{A}^{p}$ to be non-increasing, which holds if $v(t,p_{t},Z_{t})\geq 0$ for all $p\in\mathcal{A}$, and there exists a $p^{*}\in\mathcal{A}$ such that $v(t,p^{*}_{t},Z_{t})=0$. Therefore, if we choose
@@
Thus, the family of processes $\left\{V^{(p)}(\cdot,x):p\in\mathcal{A}\right\}$ becomes a collection of supermartingales. Moreover, for $p_{t}^{*}\in\Pi_{C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)}\left(Z_{t}+\frac{\theta_{t}}{\alpha}\right)$, we have $v(\cdot,p^{*},Z)=0$ for every $t\in[0,T]$, hence $V^{(p^{*})}(\cdot,x)$ becomes a true martingale.
We have now completed the construction, and it is clear how the utility maximization problem in our setting connects to a class of mean-field quadratic BSDEs (cf. Hibon et al. (2023)). In Hu et al. (2005), the optimization problem is solved under the assumptions that the random endowment $F$ is bounded and that the trading strategy $p$ takes values in an admissible set $\mathcal{A}$. However, their admissible set differs from ours: in their case, the constraint set for trading strategies is simply closed in $\mathbb{R}^{d}$. In contrast, we introduce a constraint set in which the trader’s behavior depends on the market (i.e., the behavior of other traders) in a mean-field sense. In particular, the driver depends on the expectation of the control process $Z$, and the constraint set becomes time-dependent, reflecting the market’s influence at each moment. To connect the control process $Z$ in BSDE (6.31) with the trading strategy $p\in\mathcal{A}$, we provide the following remarks bridging the mathematical framework with financial interpretations.
@@
Thus, the process $Z_{t}=p_{t}^{*}-\frac{\theta_{t}}{\alpha}$ can be interpreted as the unconstrained trading strategy adjusted by $\frac{\theta_{t}}{\alpha}$ at time $t\in[0,T]$. In general, the constraint set $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)$ depends on the average of the adjusted trading strategies among the market participants. The convergence is guaranteed by the homogeneity and exchangeability of market participants.